Johnny and Wendy were walking down the street when they saw a mysterious man frozen in ice, just like the one below. Johnny said the man should be brought to the police station and started to push the man to the east with 50 N. Wendy, on the other hand, said the man should be taken to a hospital and started pushing the man to the west with 40 N. Given that the man was inside the block of ice, assume there was no frictional force between the man and ice and the ground. The mass of the man and the ice was 75 kg. 8.

media%2Ffec%2Ffecd14b7-4a06-4c4c-9cb7-72
Answer #1

Net tonet and ae eele7aおon of the mant bloek wil be dsas Net oace om te mal i 50-1 SON 40N es Free body diagnam (FBD) Accelesnationmon in ite, at 0 uwe have, bistamer travelled by Ahelox, s-001 = 0 +0.07X3 =0.63m Let, v be the imas velaeit